OpenStudy (milliedelongg):

eliminate 6x=18-12 y and 7x+4y=9 need both x and y individually I really don't understand how to do it.

OpenStudy (campbell_st):

they are asking you to find the point of intersection for the 2 linear equations... it looks like you need to use the elimination method so take the 1st equation and divide every term by 6... as 6 is a common factor you get x = 3 - 2y for me, the easy method is to use substitution... so you take the equation above and substitute it into 7x + 4y = 9 or 7(3 - 2y) + 4y = 9 distribute and solve for y can you check the equation about and make sure it is = 9 and not = -9 when you find y, substitute it into either equation to find x. hope it helps

OpenStudy (volpina):

you could also graph the two lines and look at the point they intersect.

OpenStudy (wolf1728):

Or you can try the elimination method: 6x=18-12 y 7x+4y=9 Let's rearrange the 1st equation a) 6x +12y=18 b) 7x+4y=9 Multiply equation "b" by -3 a) 6x +12y=18 b) -21x -12y=-27 Adding both equations: -15x = -9 x = .6 and now you can easily solve for y

OpenStudy (mathmale):

"eliminate 6x=18-12 y and 7x+4y=9 need both x and y individually" would be better written as "Solve the system of linear equations 6x=18-12y and 7x+4y=9 for x and y through elimination." In other words, eliminate one variable (by addition / subtraction or substitution, find the value of the other variable, and then use one of the above equations to find the 2nd unknown variable.